OpenStudy (anonymous):

Determine the point at which the function r(t)= has a tangent line that is perpendicular to the plane 4x+y-2x=12

OpenStudy (amistre64):

what is our r' ?

OpenStudy (anonymous):

Wouldn't it be t^2, t, and 2t?

OpenStudy (amistre64):

y'=4 but yeah

OpenStudy (amistre64):

now what do we know about perp vectors and a dot product?

OpenStudy (amistre64):

over thought that ... the normal to the plane is already perp to it, so we just equate it to the normal

OpenStudy (anonymous):

Yeah I was about to say, don't they intersect?

OpenStudy (amistre64):

<t^2, 4, 2t> = n<4,1,-2> assuming thats a -2z for the plane

OpenStudy (anonymous):

Okay so now that we have that can we take the dot product?

OpenStudy (amistre64):

dot product would give us a vector parallel to the plane

OpenStudy (anonymous):

So cross?

OpenStudy (amistre64):

as is, perp to the plane is in the same direction as the normal recall that given a vector (a,b,c) and all vectors (x,y,z) perp to it; then the dot product gives us the equation of the plane ax + by + cz = 0

OpenStudy (amistre64):

so if we equate this to some scaled version of (a,b,c) then the line is in the same direction as the normal vector and is perp to the plane

OpenStudy (anonymous):

<4,1,-2> could be our (a,b,c) ?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

And we can put that into our plane?

OpenStudy (anonymous):

Or use the points from the plane as our (x,y,z)?

OpenStudy (amistre64):

no, if we equate our r' parts with a scaled version of our (a,b,c) parts, then we will find a t value tha twill define the point we want

OpenStudy (amistre64):

we want the point (t^2,4,2t) to equal some scaled version of (4,1,-2) since we need y=4, let n=4 t^2,4,2t = 4(4),1(4),-2(4) therefore t^2 = 16 and 2t = -8 and 4=4

OpenStudy (amistre64):

solve for t

OpenStudy (anonymous):

t=1?

OpenStudy (amistre64):

1^2 not= 16 -2(1) not = -8 so im going with no, its not t=1

OpenStudy (amistre64):

2(1) not = -8 .... either way lol